OpenStudy (anonymous):

The graph of y = x^2 is shown below. Using complete sentences, explain how the graph of y = -2(x + 3)^2 + 1 would differ from this parabola?

OpenStudy (anonymous):

@Callisto @Compassionate @iambatman @phi @PRAETORIAN.10 @texaschic101 @thomaster @Zarkon Can you help me?

OpenStudy (anonymous):

khbjhgjiehbgi

OpenStudy (anonymous):

@iGreen @ganeshie8

OpenStudy (igreen):

Use https://www.desmos.com/calculator to graph it, then observe the differences.

OpenStudy (igreen):

After graphing I have observed: Parabola y = x^2 is wider and opens upward. It's vertex is at the origin(0, 0). Parabola y = -2(x + 3)^2 + 1 is thinner than parabola y = x^2 and opens downward. It's vertex is at (-3, 1).

OpenStudy (anonymous):

Same! Thank you! @iGreen